OpenStudy (anonymous):

1.A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain 2.A highway makes an angle of 6 with the horizontal. This angle is maintained for a horizontal distance of 5 miles. To the nearest hundredth of a mile, how high does the highway rise in this 5-mile section? Show the steps you use to find the distance. 3.A forest ranger spots a fire from a 28-foot tower. The angle of depression from the tower to the fire is 11 . To the nearest foot, how far is the fire from the base of the tower? Show the steps you use to find the solution.

OpenStudy (raden):

for no.1 which one is true, of 15^2 = 10^2 + 7^2 or 15^2 > 10^2 + 7^2 or 15^2 < 10^2 + 7^2 ??? * if 15^2 = 10^2 + 7^2, then it called a right triangle ** if 15^2 > 10^2 + 7^2, then it called an abtuse triangle *** if 15^2 < 10^2 + 7^2, then it called a cute triangle

OpenStudy (raden):

* obtuse